OpenStudy (anonymous):

Help? Maybe? Idk? 4x+3y+82ab(4x(12y+77a)^4)-11a+2b

OpenStudy (anonymous):

@PaulaLovesSchool13

OpenStudy (anonymous):

WHAT THE HECK

OpenStudy (igreen):

\(4x+3y+82ab(4x(12y+77a)^4)-11a+2b\) Distribute the exponent 4 into the parenthesis. Remember \((a + b)^x = a^x + b^x\)

OpenStudy (anonymous):

what you have never done this stuff before? @PaulaLovesSchool13

OpenStudy (anonymous):

Uh...let the mathlete help you out for now

OpenStudy (anonymous):

I am in 8th grade dude

OpenStudy (igreen):

I have it's very easy.

OpenStudy (anonymous):

What grade are you in @iGreen

OpenStudy (igreen):

10th, but I didn't learn this in 10th..I learned in like 8th or 9th.

OpenStudy (anonymous):

I didn't learn this yet o_O

OpenStudy (anonymous):

its distributing exponents and adding like terms

OpenStudy (anonymous):

o_O well good luck

OpenStudy (anonymous):

I might stick in here in case I learn this o_O

OpenStudy (anonymous):

just randomly made this for an experiment and wow the results are stunning

OpenStudy (anonymous):

so many people came to my question when i put a few key words in the title

OpenStudy (anonymous):

Bahahahhahahah ^

OpenStudy (anonymous):

i took something most people would find simple and made it complicated by adding an exponent and extending it

OpenStudy (igreen):

\(4x+3y+82ab(4x(12y+77a)^4)−11a+2b\) Distribute the exponent 4 into the parenthesis. Remember \((a+b)^x=a^x+b^x\) Given that: \((12y + 77a)^4 = 12y^4 + 77a^4\) Put that back in the expression: \(4x+3y+82ab(4x(12y^4 + 77a^4))−11a+2b\) Now distribute the 4x into the parenthesis: \(4x(12y^4 + 77a^4) = (4x \cdot 12y^4) + (4x \cdot 77a^4)\) Given that: \(48xy^4 + 308ax^4\) Put it back into the expression: \(4x+3y+82ab(48xy^4 + 308ax^4)−11a+2b\) Distribute 82ab into the parenthesis: \(82ab(48xy^4 + 308ax^4) = 3936abxy^4 + 25256a^2bx^4\) Put it back in the expression: \(4x+3y+3936abxy^4 + 25256a^2bx^4−11a+2b\) Reorder the terms: \(25256a^2bx^4 + 3936abxy^4 - 11a + 4x + 3y + 2b\)

OpenStudy (anonymous):

omg

OpenStudy (anonymous):

A medal for you, for sure

OpenStudy (anonymous):

lol i didnt feel like answering that but he did a great job @iGreen you should take pre-calculas in your senior year

OpenStudy (anonymous):

bahahahha ^

OpenStudy (igreen):

Lol. That's not pre-calculus it's too easy.. xD

OpenStudy (anonymous):

Easy? I wish...

OpenStudy (anonymous):

i know its not but you might like pre-calculus